# Rabi oscillations and entanglement between two atoms interacting by the Rydberg blockade and with a quantized radiation field studied by the Jaynes-Cummings Model

### v1

@@ 4 Coherent states of the field
Figure 7 shows the atoms-field entanglement $\varepsilon(t)$ in (66) for zero detuning ($\Delta=0$) and $\bar{n}=10$, $20$, and $50$ as a function of $t$ (orange lines). As expected, $\varepsilon(0)=0$ for the initial state |$\ket{\psi^{(\alpha)}(0)}$ in (56); subsequently, $\varepsilon(t)$ increases to its maximum value $1/2$ as $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ (blue lines in Figure 7) collapses to a constant value $1/2$. Interestingly, $\varepsilon(t)$ is not constant but decreases to a value close to 0 during the period where $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ remains constant at $\simeq 1/2$ after the initial collapse, indicating a nontrivial correlation between the probability $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ and the atoms-field entanglement $\varepsilon(t)$.
5 Summary and outlook
@@
Finally, we find explicit expressions for the dynamics of the combined system (atoms + field) for the case that the field is in a quantum single-mode coherent state, and discuss the entanglement between the two-atom system and the coherent field in terms of the parameter $\varepsilon(t)$ in (66) (Section 4 and Figure 7). Interestingly, $\varepsilon(t)$ is not constant but decreases to a value close to 0 during an extended period in which the probability of the atomic state |$\ket{\text{sym}}=\tfrac{1}{\sqrt{2}}\left(\ket{r,g}+\ket{g,r}\right)$ remains constant at the value $1/2$ (Figure 7). The non-trivial correlation between the atomic states and the atom-field entanglement $\varepsilon(t)$ is potentially relevant for quantum state-transfer between atoms and photons in quantum information processing.
@@
6 Acknowledgements
@@
In this appendix we first derive results for the initial state |$\ket{\psi_{B}^{(n)}(0)}=\ket{\psi_{3}^{(n)}}=\ket{g,g,n+1}$ discussed in Section 3.2 (Appendix A.1), and then for the complementary initial state |$\ket{\psi_{\beta}^{(n)}(0)}:=\ket{\text{sym},n}=\frac{1}{\sqrt{2}}\left(\ket{\psi_{1}^{(n)}}+\ket{\psi_{2}^{(n)}}\right)$ (Appendix A.2).
@@
$=\frac{1}{\sqrt{2}}\,\frac{1}{2}\sin(2\varphi_{n})\exp\Big{[}-i\Big{(}\omega_{f}n+\tfrac{\Delta}{2}\Big{)}t\Big{]}\left[\exp\left(-i\Omega_{n}t\right)-\exp\left(i\Omega_{n}t\right)\right]$ (76)
$=\frac{1}{\sqrt{2}}(-i)\sin(2\varphi_{n})\exp\Big{[}-i\Big{(}\omega_{f}n+\tfrac{\Delta}{2}\Big{)}t\Big{]}\sin(\Omega_{n}t)\,,$ (77)
@@
$=\exp\Big{[}-i\Big{(}\omega_{f}n+\tfrac{\Delta}{2}\Big{)}t\Big{]}\left[\sin^{2}(\varphi_{n})\exp\left(-i\Omega_{n}t\right)+\cos^{2}(\varphi_{n})\exp\left(i\Omega_{n}t\right)\right]\,.$ (82)
@@
For later use (Appendix B), here we quote results for the coefficients $\mu_{\beta}^{(n)}(t)$, $\nu_{\beta}^{(n)}(t)$, $\xi_{\beta}^{(n)}(t)$ according to (36) and the resulting probabilities for the initial state |$\ket{\psi_{\beta}^{(n)}(0)}:=\ket{\text{sym},n}$, which by (46) and (87) is complementary to the initial state |$\ket{\psi_{B}^{(n)}(0)}=\ket{\psi_{3}^{(n)}}$ in Appendix A.1. Similarly as in Appendix A.1 one finds
@@
$\mu_{\beta}^{(n)}(t)$ $=\left\langle\psi_{1}^{(n)}\Big{|}\psi_{\beta}^{(n)}(t)\right\rangle$ (91)
$=\frac{1}{\sqrt{2}}\exp\Big{[}-i\Big{(}\omega_{f}n+\tfrac{\Delta}{2}\Big{)}t\Big{]}\left[\cos^{2}(\varphi_{n})\exp\left(-i\Omega_{n}t\right)+\sin^{2}(\varphi_{n})\exp\left(i\Omega_{n}t\right)\right]\,,$ (92)
@@
$\xi_{\beta}^{(n)}(t)$ $=\left\langle\psi_{3}^{(n)}\Big{|}\psi_{\beta}^{(n)}(t)\right\rangle$ (94)
$=(-i)\sin(2\varphi_{n})\exp\Big{[}-i\Big{(}\omega_{f}n+\tfrac{\Delta}{2}\Big{)}t\Big{]}\sin(\Omega_{n}t)\,.$ (95)
@@
For the initial states |$\ket{\psi_{C}^{(n)}(0)}=\ket{\psi_{1}^{(n)}}=\ket{r,g,n}$ considered in Section 3.3 and |$\ket{\psi_{2}^{(n)}}=\ket{g,r,n}$ the time evolution |$\ket{\psi^{(n)}(t)}$ can be deduced from case A (Section 3.1) and Appendix A.2, using
@@
with |$\ket{\psi_{\beta}^{(n)}(0)}=\frac{1}{\sqrt{2}}\left(\ket{\psi_{1}^{(n)}}+\ket{\psi_{2}^{(n)}}\right)$ (Appendix A.2) and |$\ket{\psi_{A}^{(n)}(0)}=\frac{1}{\sqrt{2}}\left(\ket{\psi_{1}^{(n)}}-\ket{\psi_{2}^{(n)}}\right)$ (Section 3.1). For the initial state |$\ket{\psi_{C}^{(n)}(0)}=\ket{\psi_{1}^{(n)}}$, (98) implies for the coefficients in (36)
@@
Appendix C Derivation of $\gamma(t)$ and $\left|\gamma(t)\right|^{2}$ in (68), (69)
$\gamma(t)$ $=\langle\text{sym}\,|\,\hat{\rho}_{\text{atoms}}(t)\,|\,g,g\rangle=\langle\text{sym}\,|\,\text{tr}_{f}\left\{\hat{\rho}(t)\right\}|\,g,g\rangle$ (103)
@@
Using (67) and the normalization condition $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)+P_{\text{atoms}}^{(\alpha)}\left(\ket{g,g},t\right)=1$ one obtains
$\rho_{\text{atoms}}^{2}(t)=\begin{pmatrix}P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)^{2}+\left|\gamma(t)\right|^{2}&\gamma(t)\\[4.0pt] \gamma^{*}(t)&P_{\text{atoms}}^{(\alpha)}\left(\ket{g,g},t\right)^{2}+\left|\gamma(t)\right|^{2}\end{pmatrix}$ (109)
which yields (69). For $\left|\gamma(t)\right|^{2}$ we only quote the result for $\Delta=0$ (as used in Figure 7) for simplicity, using (107):
@@
$\left|\gamma(t)\right|^{2}=\exp\left(-2\bar{n}\right)\bar{n}\left[\,\sum\limits_{k=0}^{\infty}\frac{\bar{n}^{k}}{k!}\frac{1}{\sqrt{k+1}}\sin\left(\lambda\sqrt{2(k+1}\,t\right)\cos\left(\lambda\sqrt{2k}\,t\right)\right]^{2}\,,\,\,\Delta=0\,,$ (110)
where $\bar{n}=\left|\alpha\right|^{2}$.

### v2

@@ 5.3 Entanglement of formation $E(a-b)$ between the two atoms $a$ and $b$
Figure 8 shows that the three forms of entanglement $E(ab-f,t)$, $E(a-bf,t)$, $E(a-b;t)$ (Figure 7) and the probability $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ are strongly correlated. The direct correlation between $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ and $E(a-b;t)$ is evident by (78) and Figure 9. As expected, all three forms of entanglement vanish for $t=0$ because the total system is initially prepared in the product state |$\ket{g,g,\alpha}$ in (57). Subsequently, $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ collapses to a constant value $1/2$ (orange lines in parts (A), (C), (E)). Interestingly, $E(ab-f,t)$ and $E(a-bf,t)$ (blue and magenta lines in parts (B), (D), (F)) do not remain constant but decrease to relatively low values during the period where $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ remains constant at $\simeq 1/2$ after the initial collapse.
6 Summary and conclusions
@@
We find explicit expressions for the dynamics of the combined system (atoms + field) for the case that the field is in a quantum single-mode coherent state (Section 4 and Figure 8). Using these results, we obtain the entanglement between various components of the total system consisting of the two Rydberg-interacting atoms $a$, $b$ and the quantized radiation field (Figures 7 and 8). We find distinct correlations between these entanglements and the probability $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ of finding the atoms in the symmetric state |$\ket{\text{sym}}=\frac{1}{\sqrt{2}}\left(\ket{r,g}+\ket{g,r}\right)$ in (60), which are absent for a system of non-interacting atoms studied by the Tavis-Cummings model (see below). The correlations between the atomic states and the atom-field entanglement is potentially relevant for quantum state-transfer between atoms and photons in quantum information processing.
It is interesting to compare the three types of entanglement for the system with Rydberg-interacting atoms shown in parts (B), (D), and (F) of Figure 8 with the corresponding entanglements for a system with non-interacting atoms described by the two-atom Tavis-Cummings model (TC). For zero detuning and identical coupling constant $\lambda$ between each of the atoms $a$ and $b$ and the field, the TC model is governed by the Hamiltonian
$\hat{H}_{TC}=\hbar\omega_{f}{\hat{a}}^{{\dagger}}{\hat{a}}-\frac{1}{2}\hbar\omega_{f}\left({\hat{\sigma}}_{z}^{a}+{\hat{\sigma}}_{z}^{b}\right)\,+\hbar\lambda\left({\hat{\sigma}}_{+}^{a}{\hat{a}}+{\hat{\sigma}}_{-}^{a}{\hat{a}}^{{\dagger}}+{\hat{\sigma}}_{+}^{b}{\hat{a}}+{\hat{\sigma}}_{-}^{b}{\hat{a}}^{{\dagger}}\right)\,,$ (79)
where $\omega_{f}$ is again the angular frequency of the quantized radiation field, which for zero detuning is equal to the atomic transition frequency $\omega_{0}$. The TC Hamiltonian (79) is obtained from the Hamiltonian (4) by replacing the projection operators ${\hat{P}}_{g}^{\,a}=\mathinner{|{g_{a}}\rangle\langle{g_{a}}|}$ and ${\hat{P}}_{g}^{\,b}=\mathinner{|{g_{b}}\rangle\langle{g_{b}}|}$ to the ground states of the atoms with the identity operator $\hat{I}$, so that the doubly excited state |$\ket{r,r,\phi_{f}}$ is accessible in the TC model. Entanglements between the atoms and the field in the TC model were studied in reference [28] in terms of tangles, which is a different (albeit closely related, see D) entanglement measure than the von Neumann entropy and entanglement of formation discussed in Section 5 and shown in our Figure 8.
To facilitate a one-to-one comparison of the behavior of the system with Rydberg-interacting atoms (this work) and independent atoms (TC model), Figure 10 shows the atomic probabilities $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$, $P_{\text{atoms}}^{(\alpha)}\left(\ket{gg},t\right)$, and $P_{\text{atoms}}^{(\alpha)}\left(\ket{rr},t\right)$ of the respective atomic states |$\ket{\text{sym}}$, |$\ket{gg}$, and |$\ket{rr}$ (see (126) in D), as well as the three types of entanglement $E(ab-f;t)$, $E(a-bf;t)$, and $E(a-b;t)$ defined in (70), (73), and (78), respectively, obtained for the TC model in (79) (see D for details). Figure 10 should be compared with Figure 8. The parameters, labels, and the meaning of the colored lines in Figure 10 are the same as in Figure 8; the only exception are the probabilities $P_{\text{atoms}}^{(\alpha)}\left(\ket{rr},t\right)$ shown as purple lines in parts (A), (C), (E) in Figure 10, which are absent in Figure 8 because the doubly excited state |$\ket{rr}$ is not accessible for Rydberg-interacting atoms.
Figures 8 and 10 show that the time dependence of the three types of entanglement in parts (B), (D), and (F) is notably different for Rydberg-interacting atoms compared to the TC model. In particular, for the TC model (Figure 10), the entanglements are less correlated with the atomic probability $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ than for the system with Rydberg-interacting atoms (Figure 8). Moreover, the atom-atom entanglement $E(a-b;t)$ is significantly lower for the TC model than for the system with Rydberg-interacting atoms. Thus, the Rydberg blockade interaction between the atoms strongly affects entanglement in the system overall and increases the entanglement between the atoms.
As mentioned above, in reference [28], entanglements between the atoms and the field in the TC model were studied in terms of tangles instead of the von Neumann entropy and entanglement of formation shown in Figures 8 and 10. In Figure 11 in D we show tangles corresponding to the system in Figure 10 for $\bar{n}=50$ and $\bar{n}=100$. The curves for $\bar{n}=50$ in Figure 11A closely resemble the corresponding curves in Figure 10F, showing that our conclusions are independent of the particular entanglement measure. The curves for $\bar{n}=100$ in Figure 11B exactly reproduce the corresponding curves in Fig. 2a in [28].
@@
Figure 10: Time-dependence, in units of $\lambda^{-1}$, of the atomic probabilities $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$, $P_{\text{atoms}}^{(\alpha)}\left(\ket{gg},t\right)$, and $P_{\text{atoms}}^{(\alpha)}\left(\ket{rr},t\right)$ (see (126) in D), as well as the three types of entanglement $E(ab-f;t)$, $E(a-bf;t)$, and $E(a-b;t)$ defined in (70), (73), and (78), respectively, obtained for the TC model in (79) (compare Figure 8). The sum of the atomic probabilities is equal to 1 by normalization. The parameters, labels and the meaning of the colored lines are the same as in Figure 8; the only exception are the probabilities $P_{\text{atoms}}^{(\alpha)}\left(\ket{rr},t\right)$ shown as purple lines in parts (A), (C), (E), which are absent in Figure 8 because the doubly excited state |$\ket{rr}$ is not accessible for Rydberg-interacting atoms. As in Figure 8, results are shown for $\bar{n}=10$, $20$, and $50$ for $\omega_{f}=\omega_{0}=\lambda$, zero detuning ($\Delta=0$), and for the initial state |$\ket{g,g,\alpha}$. See text and D.
Acknowledgements
@@
We first derive results for the initial state |$\ket{\psi_{B}^{(n)}(0)}=\ket{\psi_{3}^{(n)}}=\ket{g,g,n+1}$ discussed in Section 3.2 (A.1), and then for the complementary initial state |$\ket{\psi_{\beta}^{(n)}(0)}:=\ket{\text{sym},n}=\frac{1}{\sqrt{2}}\left(\ket{\psi_{1}^{(n)}}+\ket{\psi_{2}^{(n)}}\right)$ (A.2).
@@
$=\frac{1}{\sqrt{2}}\,\frac{1}{2}\sin(2\varphi_{n})\exp\Big{[}-i\hskip 1.0pt\Big{(}\omega_{f}n+\tfrac{\Delta}{2}\Big{)}\hskip 1.0ptt\Big{]}\left[\exp\left(-i\Omega_{n}t\right)-\exp\left(i\Omega_{n}t\right)\right]$ (86)
[…]
@@
$=\exp\Big{[}-i\hskip 1.0pt\Big{(}\omega_{f}n+\tfrac{\Delta}{2}\Big{)}\hskip 1.0ptt\Big{]}\left[\sin^{2}(\varphi_{n})\exp\left(-i\Omega_{n}t\right)+\cos^{2}(\varphi_{n})\exp\left(i\Omega_{n}t\right)\right]\,.$ (92)
@@
For later use (B), here we quote results for the coefficients $\mu_{\beta}^{(n)}(t)$, $\nu_{\beta}^{(n)}(t)$, $\xi_{\beta}^{(n)}(t)$ according to (37) and the resulting probabilities for the initial state |$\ket{\psi_{\beta}^{(n)}(0)}:=\ket{\text{sym},n}$, which by (47) and (97) is complementary to the initial state |$\ket{\psi_{B}^{(n)}(0)}=\ket{\psi_{3}^{(n)}}$ in A.1. Similarly as in A.1 one finds
@@
$\mu_{\beta}^{(n)}(t)$ $=\left\langle\psi_{1}^{(n)}\Big{|}\psi_{\beta}^{(n)}(t)\right\rangle=\frac{1}{\sqrt{2}}\exp\Big{[}-i\hskip 1.0pt\Big{(}\omega_{f}n+\tfrac{\Delta}{2}\Big{)}\hskip 1.0ptt\Big{]}\left[\cos^{2}(\varphi_{n})\exp\left(-i\Omega_{n}t\right)+\sin^{2}(\varphi_{n})\exp\left(i\Omega_{n}t\right)\right]\,,$ (101)
@@
$\xi_{\beta}^{(n)}(t)$ $=\left\langle\psi_{3}^{(n)}\Big{|}\psi_{\beta}^{(n)}(t)\right\rangle=(-i)\sin(2\varphi_{n})\exp\Big{[}-i\hskip 1.0pt\Big{(}\omega_{f}n+\tfrac{\Delta}{2}\Big{)}\hskip 1.0ptt\Big{]}\sin(\Omega_{n}t)\,.$ (103)
@@
For the initial states |$\ket{\psi_{C}^{(n)}(0)}=\ket{\psi_{1}^{(n)}}=\ket{r,g,n}$ considered in Section 3.3 and |$\ket{\psi_{2}^{(n)}}=\ket{g,r,n}$ the time evolution |$\ket{\psi^{(n)}(t)}$ can be deduced from case A (Section 3.1) and A.2, using
@@
with |$\ket{\psi_{\beta}^{(n)}(0)}=\frac{1}{\sqrt{2}}\left(\ket{\psi_{1}^{(n)}}+\ket{\psi_{2}^{(n)}}\right)$ (A.2) and |$\ket{\psi_{A}^{(n)}(0)}=\frac{1}{\sqrt{2}}\left(\ket{\psi_{1}^{(n)}}-\ket{\psi_{2}^{(n)}}\right)$ (Section 3.1). For the initial state |$\ket{\psi_{C}^{(n)}(0)}=\ket{\psi_{1}^{(n)}}$, (106) implies for the coefficients in (37)
@@
Appendix C Derivation of $\gamma(t)$ in (72)
$\gamma(t)$ $=\langle\text{sym}\,|\,\hat{\rho}_{\text{atoms}}(t)\,|\,g,g\rangle=\langle\text{sym}\,|\,\text{tr}_{f}\left\{\hskip 1.0pt\hat{\rho}(t)\right\}|\,g,g\rangle$ (111)
@@ Appendix C Derivation of $\gamma(t)$ in (72)
Appendix D Tavis-Cummings model
Here we briefly review results for the two-atom Tavis-Cummings (TC) model described by the Hamiltonian (79) as needed to derive the probabilities and entanglements for the TC model shown in Figure 10, according to the notation and definitions used in Sections 2 - 5. These results can be deduced from previous work, see, e.g., references [28, 81].
The total Hilbert space $\mathbb{H}$ of the TC model (79) decouples into 4-dimensional subspaces
$\mathbb{H}_{n}=\text{span}\left\{\ket{r,r,n-1},\ket{r,g,n},\ket{g,r,n},\ket{g,g,n+1}\right\}\,,\,\,n=1,2,\ldots$ (117)
including the doubly excited atomic state |$\ket{r,r,n-1}$ (compare the text in the second paragraph below equation (1) in [28]). As in Section 2, the index $n$ for $\mathbb{H}_{n}$ corresponds to the number of photons in the states |$\ket{r,g,n}$ and |$\ket{g,r,n}$ (see text below (11)); for $n=0$, the Hilbert space is spanned by the 3 states |$\ket{r,g,0}$, |$\ket{g,r,0}$, |$\ket{g,g,1}$, and for the index $n=-1$ only the state |$\ket{g,g,0}$ is available to the system. In the basis corresponding to (117) the Hamiltonian ${\hat{H}}_{TC}^{\,(n)}$ is represented by a $4\times 4$ matrix
${\hat{H}}_{TC}^{\,(n)}=\hbar\begin{pmatrix}n\omega_{f}&\lambda\sqrt{n}&\lambda\sqrt{n}&0\\[4.0pt] \lambda\sqrt{n}&n\omega_{f}&0&\lambda\sqrt{n+1}\\[4.0pt] \lambda\sqrt{n}&0&n\omega_{f}&\lambda\sqrt{n+1}\\[4.0pt] 0&\lambda\sqrt{n+1}&\lambda\sqrt{n+1}&n\omega_{f}\end{pmatrix}\,\,,\,n=1,2,\ldots$ (118)
For the initial state |$\ket{\psi^{(n)}_{TC}(0)}=\ket{g,g,n+1}$ at $t=0$ (corresponding to case B discussed in Section 3.2) the time evolution of the system is given by
|$\ket{\psi^{(n)}_{TC}(t)}$ $=\exp\left(-\frac{i}{\hbar}\hat{H}_{TC\,}t\right)\ket{g,g,n+1}$
$=\rho_{n}(t)\ket{r,r,n-1}+\sigma_{n}(t)\ket{\text{sym},n}+\xi_{n}(t)\ket{g,g,n+1}\,\,,\,n=1,2,\ldots$ (119)
where
$\rho_{n}(t)$ $=\exp(-i\omega_{f}nt)\,a(n)\,b(n)\left(\cos\left(\Omega_{n}^{TC}t\right)-1\right)\,,$ (120)
$\sigma_{n}(t)$ $=-i\exp(-i\omega_{f}nt)\,a(n)\sin\left(\Omega_{n}^{TC}t\right)\,,$ (121)
$\xi_{n}(t)$ $=\exp(-i\omega_{f}nt)\left(a(n)^{2}\cos\left(\Omega_{n}^{TC}t\right)+b(n)^{2}\right)\,,$ (122)
with $a(n)=\sqrt{(n+1)/(2n+1)}$, $b(n)=\sqrt{n/(2n+1)}$, and the Rabi frequency of the Tavis-Cummings model  [^10] Note that $\Omega_{n}^{TC}$ in (123) is different from the Rabi frequency $\Omega_{n}(0)$ in (30) by an extra factor of 2 in front of $n$.
$\Omega_{n}^{TC}=\lambda\sqrt{2(2n+1)}\,\,.$ (123)
The normalization $\left|\hskip 1.0pt\rho_{n}(t)\right|^{2}+\left|\hskip 1.0pt\sigma_{n}(t)\right|^{2}+\left|\hskip 1.0pt\xi_{n}(t)\right|^{2}=1$ is fulfilled due to $a(n)^{2}+b(n)^{2}=1$.
If the two atoms are initially in their ground states |$\ket{g}$ and the field is in a quantum single-mode coherent state |$\ket{\alpha}$ as in (56), the initial state |$\ket{\psi_{TC}^{(\alpha)}(0)}=\ket{g,g,\alpha}$ evolves to
@@ Appendix D Tavis-Cummings model
|$\ket{\psi_{TC}^{(\alpha)}(t)}=\exp\left(-\frac{\left|\alpha\right|^{2}}{2}\right)\left[\exp(i\omega_{0}t)\ket{g,g,0}+\sum\limits_{n=0}^{\infty}\frac{\alpha^{n+1}}{\sqrt{(n+1)!}}\ket{\psi^{(n)}_{TC}(t)}\right]\,,\,\,t>0\,,$ (124)
with |$\ket{\psi^{(n)}_{TC}(t)}$ in (D) (compare (59)). The reduced density operator for the atoms, $\hat{\rho}^{TC}_{\text{atoms}}(t)$, is given by the partial trace of the density operator $\hat{\rho}_{TC}(t)$ of the total system (atoms + field) in the pure state |$\ket{\psi_{TC}^{(\alpha)}(t)}$ over the field |$\ket{\phi_{f}}$, i.e.,
$\hat{\rho}^{TC}_{\text{atoms}}(t)=\text{tr}_{f}\left\{\hskip 1.0pt\hat{\rho}_{TC}(t)\right\}=\sum\limits_{k=0}^{\infty}\langle k\,|\,\psi_{TC}^{(\alpha)}(t)\rangle\langle\psi_{TC}^{(\alpha)}(t)\,|\,k\rangle\,,$ (125)
where |$\ket{k}$, $k=0,1,2,\ldots$ (compare (62) and (63)). In the basis {|$\left\{\ket{rr},\ket{\text{sym}},\ket{gg}\right\}$, with |$\ket{\text{sym}}=\frac{1}{\sqrt{2}}\left(\ket{r,g}+\ket{g,r}\right)$, the reduced density operator for the atoms is represented by the matrix
$\rho^{TC}_{\text{atoms}}(t)=\begin{pmatrix}P_{\text{atoms}}^{(\alpha)}\left(\ket{rr},t\right)&\theta^{*}(t)&\phi(t)\\[4.0pt] \theta(t)&P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)&\gamma(t)\\[4.0pt] \phi^{*}(t)&\gamma^{*}(t)&P_{\text{atoms}}^{(\alpha)}\left(\ket{gg},t\right)\hskip 1.0pt\end{pmatrix}\,\,.$ (126)
[…]
$\overline{\rho}^{TC}_{\text{atoms}}(t)=\begin{pmatrix}P_{gg}(t)&\frac{1}{\sqrt{2}}\gamma^{*}(t)&\frac{1}{\sqrt{2}}\gamma^{*}(t)&\phi^{*}(t)\\[4.0pt] \frac{1}{\sqrt{2}}\gamma(t)&\frac{1}{2}P_{\text{sym}}(t)&\frac{1}{2}P_{\text{sym}}(t)&\frac{1}{\sqrt{2}}\theta(t)\\[4.0pt] \frac{1}{\sqrt{2}}\gamma(t)&\frac{1}{2}P_{\text{sym}}(t)&\frac{1}{2}P_{\text{sym}}(t)&\frac{1}{\sqrt{2}}\theta(t)\\[4.0pt] \phi(t)&\frac{1}{\sqrt{2}}\theta^{*}(t)&\frac{1}{\sqrt{2}}\theta^{*}(t)&P_{rr}(t)\end{pmatrix}\,\,,$ (127)
where the overbar indicates that the matrix is with respect to the standard basis. We use the same type of abbreviation for the probabilities as quoted below (77) with $P_{gg}(t)+P_{\text{sym}}(t)+P_{rr}(t)=1$ by normalization. The three types of entanglement $E(ab-f;t)$, $E(a-bf;t)$, and $E(a-b;t)$ shown in Figure 10 for the TC model were calculated using the density matrices $\rho^{TC}_{\text{atoms}}(t)$ and $\overline{\rho}^{TC}_{\text{atoms}}(t)$ in (126) and (127), respectively, as described in Section 5.
In reference [28], entanglements between the atoms and the field in the TC model were studied in terms of tangles, which is a different entanglement measure than the von Neumann entropy and entanglement of formation discussed in Section 5 and shown in our Figures 8 and 10 (see equations (7), (8), and (2) in [28]). In terms of tangles, the entanglement between the two-atom system $ab$ and the field $f$ is given by (instead of $E(ab-f;t)$ in (70))
$\tau(ab-f;t)=2\left[1-\text{tr}\left\{\left(\overline{\rho}^{TC}_{\text{atoms}}(t)\right)^{2}\right\}\right]\,.$ (128)
The entanglement between atom $a$ and the (atom $b$)-field system $bf$ is given by (instead of $E(a-bf;t)$ in (73))
$\tau(a-bf;t)=2\left[1-\text{tr}\left\{\left(\rho^{TC}_{a}(t)\right)^{2}\right\}\right]\,,$ (129)
where the reduced density matrix $\rho^{TC}_{a}$ of atom $a$ is the partial trace of $\overline{\rho}^{TC}_{\text{atoms}}(t)$ over atom $b$ (compare (74)). Finally, the atom-atom entanglement is given by the tangle (instead of $E(a-b;t)$ in (78))
$\tau(a-b;t)=C\left(\hskip 1.0pt\overline{\rho}^{TC}_{\text{atoms}}(t)\right)^{2}\,,$ (130)
where $C\left(\hskip 1.0pt\overline{\rho}^{TC}_{\text{atoms}}(t)\right)=\max\left\{0,\lambda_{1}-\lambda_{2}-\lambda_{3}-\lambda_{4}\right\}$ is the concurrence for the density matrix $\overline{\rho}^{TC}_{\text{atoms}}(t)$ (see Section 5.3; here, 3 of the 4 $\lambda_{i}$’s are nonzero). For reference purposes, Figure 11 shows the tangles in (128) - (130) for $\bar{n}=50$ and $\bar{n}=100$ for the same system as considered in Figure 10. The curves for $\bar{n}=50$ in Figure 11A closely resemble the corresponding curves in Figure 10F, and the curves for $\bar{n}=100$ in Figure 11B exactly reproduce the corresponsing curves in Fig. 2a in [28].
Figure 11: Tangles in (128) - (130) for (A) $\bar{n}=50$ and (B) $\bar{n}=100$ for the same system as considered in Figure 10.